A particle moves along a straight line OX .at a time t the distance x of the particle from O  is given by x=40 +12t-t^3.how long would the particle travel before coming to rest

Grade:11

3 Answers

Aisha
6 years ago
We have equationx=40+12t-t^3When t=0Then x=40 nowdx/dy = 12-3t^2 when velocity is zero then t=2When t=2 then x=56 now we can find how long object move before come to zero 56-40=16m
Vaidik Trivedi
5 years ago
We have equation x= 40+12t-t^3. So initially take t=0  => x=40m.
Now differtiate the equation with respect to time(t) which is equal to velocity (v) {v=dx/dt} now put v=0 as the particle is stopping you will get value of t. Put this value in "40+12t-t^3" and you will get final distance.on Subtracting the initial distance from this final distance we'll get x=16m.
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
 
x = 40 + 12t – t3
speed, v = dx/dt = 12 – 3t2
When it comes to rest, v = 0
12 – 3t2 = 0
Hence, t = 2 s
Hence, distance covered before stopping,
x = 40 + 12(2) – (2)3
   = 40 + 24 – 8
   = 56 m
 
Hope it helps.
Thanks and regards,
Kushagra